# Distributed Model Checking in Graphs Classes

### v1

@@ 2.3 Graphs of Bounded Expansion
It has been recently shown in [28] that decision, counting and optimization of FO formulas can be performed in $\mathcal{O}(\log n)$ CONGEST rounds for connected labeled graphs of bounded treedepth.
Proposition 1 ([28]).
@@
Let $\mathcal{G}^{\prime}[\Lambda]$ be a class of connected labeled graphs of bounded treedepth and let $\varphi(x_{1},\dots,x_{k})\in FO[\Lambda]$. Then, for any and any $(G,L)\in\mathcal{G}^{\prime}[\Lambda]$, the following problems can be solved by a CONGEST algorithm using $\mathcal{O}(\log n)$ rounds:
  1. 1.
@@
    Deciding $\varphi$, i.e., deciding whether it exists a $k$-uple of vertices $(v_{1},\dots,v_{k})$ such that $(G,L)\models\varphi(v_{1},\dots,v_{k})$.
  2. 2.
@@
    Counting the solutions of $\varphi$, i.e., computing the number of $k$-uples $(v_{1},\dots,v_{k})$ such that $(G,L)\models\varphi(v_{1},\dots,v_{k})$.
  3. 3.
    Optimizing $\varphi$, i.e., if graph $G$ comes with a weight function $\omega:V(G)\to\mathbb{N}$ upper bounded by some polynomial, the algorithm computes a $k$-tuple of vertices $(v_{1},\dots,v_{k})$ such that $G\models\varphi(v_{1},\dots,v_{k})$ and $\sum_{i=1}^{k}\omega(v_{i})$ is maximum (or minimum).
@@
Graphs of bounded expansion [67] can be defined in term of local minors and degeneracy as pointed out in the Introduction. It is more convenient for us to use the following characterization. Specifically, let $\mathbb{N}$ denote the set of positive integers. A class $\mathcal{G}$ of graphs has bounded expansion if there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and for every $k\in\mathbb{N}$, there exists a coloring of the vertices of $G$ with colors in $[f(k)]=\{1,\dots,f(k)\}$ such that, for every set $S\subseteq[f(k)]$ of cardinality $k$ the subgraph of $G$ induced by the nodes colored by a color in $S$ has treedepth at most $|S|$. In particular, each subset of at most $k$ colors induces a subgraph of treedepth at most $k$.
@@
For every positive integers $p$ and $k$, a $(p,k)$-treedepth coloring of a graph $G$ is a $k$-coloring of $G$ satisfying that each set $S\subseteq[k]$ of at most $p$ colors induces a graph of tree-depth at most $|S|$. Given a function $f:\mathbb{N}\rightarrow\mathbb{N}$, a class of graphs $\mathcal{G}$ has expansion $f$ if, for every $p\in\mathbb{N}_{>0}$, every $G\in\mathcal{G}$ has a $(p,f(p))$-treedepth coloring. A class of graphs has bounded expansion if it has expansion $f$ for some function $f$. A class of graphs has effective bounded expansion if it has expansion $f$ for some Turing-computable function $f$. In the following, we always assume that function $f$ is Turing-computable functions. Interestingly, treedepth colorings can be efficiently computed in the CONGEST model, as stated below.
Proposition 2 ([68]).
@@
Let $\mathcal{G}$ be a class of graphs of bounded expansion. There exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, for any graph $G\in\mathcal{G}$, and any constant positive integer $p$, a $(p,f(p))$-treedepth coloring of $G$ can be computed in $O(\log n)$ rounds in the CONGEST model.
3 Model Checking
In this section, we prove Theorem 1. We restate it here for the reader’s convenience.
See 1
@@ 3 Model Checking
We prove the theorem by adapting to our distributed setting the quantifier elimination technique of [71]. Let $(G,L)$ be a labeled graph with $G\in\mathcal{G}$, on which we aim at checking formula $\varphi$. Quantifier elimination proceeds by induction, and formula $\varphi$ may have free variables at some stages of the induction process. Our goal is to give a CONGEST algorithm that in $\mathcal{O}(D+\log n)$ transforms both the formula $\varphi(\overline{x})$ and the labeled graph $(G,L)$ into a pair $(\widehat{\varphi}(\overline{x}),\widehat{L})$ such that:
@@
  • $\widehat{\varphi}(\overline{x})$ is quantifier-free but uses new, particular unary and binary predicates.
  • $\widehat{L}$ a new set of labels for the nodes of $G$.
  • For any tuple $\overline{v}$ of $|\overline{x}|$ vertices of $G$, $(G,L)\models\varphi(\overline{v})\iff(G,\widehat{L})\models\widehat{\varphi}(\overline{v})$.
@@
Quantifiers will be eliminated one by one, and the main difficulty consists in the elimination of the existential quantifiers. As in [71], the proof works in three steps. First, the transformation from $\varphi$ and $(G,L)$ to $(\widehat{\varphi},\widehat{L})$ is performed on rooted forests of constant depth (cf. Subsection 3.1). This is actually the most technical part of the proof. Second, the quantifier elimination is performed on graphs of bounded treedepth (cf. Subsection 3.2), by reducing the analysis of graphs of bounded treedepth to the analysis of rooted forests by enriching the set of labels. Third, the analysis of graphs of bounded expansion is reduced to tha analysis of graphs of bounded treedepth thanks to $(p,f(p))$-treedepth coloring (cf. Subsection 3.3). We emphasis that, at each step of the transformation of the original formula $\varphi$, and the original labeled graph $(G,L)$, the computations must be efficiently performed in the CONGEST model (typically, one needs to make sure that adding new labels to the graph can be done efficiently in CONGEST). Luckily, the rooted forests involved in Subsection 3.1, as well as the graphs of bounded treedepth involved in Subsection 3.2, are subgraphs of the communication graph $G$, and thus all communications passing through the edges of the former can be implemented along the edges of the latter.
3.1 Rooted Forests of Constant Depth
@@
In this section we explain how to implement in CONGEST the quantifier elimination procedure from [16, 48] for rooted forests of constant depth. That is, every tree of the forest at hand is rooted, and of depth at most $d$, for some $d\geq 1$. The basic idea for eliminating the quantifiers is to encode information about the satisfiability of the formula in labels at the nodes. This increases the size of the labeling. However, by taking advantage of the simple structure of FO formulas on forests of bounded depth, it is possible to bound the number of labels by a constant depending solely on the size of the formula, and on the depth of the forests. In the following, for each positive integer $d$, we denote by $\mathcal{F}_{d}$ the set of all rooted forests of depth at most $d$. For a set of labels $\Lambda$, we denote by $\mathcal{F}_{d}[\Lambda]$ the set of all $\Lambda$-labeled graphs $(G,L)$ such that $G\in\mathcal{F}_{d}$.
3.1.1 First-Order Formulas for Forests of Bounded Depth
@@
Let $\Lambda$ be a (finite) set of labels, and let $d$ be a positive integer. It is convenient to consider a specific way of encoding formulas, suitable for $\mathcal{F}[\Lambda]$. For each $i\in\{-1,0,1,\dots,d\}$, we define the predicate $\textsf{lca}_{i}(x,y)$ as the predicate with free variables $x$ and $y$ that holds if the path $P_{x}$ from $x$ to the root of the tree containing $x$, and the path $P_{y}$ from $y$ to the root of the tree containing $y$ share exactly $i+1$ nodes. Moreover, we set $\textsf{lca}_{-1}(x,y)$ to be true whenever $x$ and $y$ are in different trees. Note that $\textsf{lca}_{i}(x,x)$ holds whenever $x$ is a node at depth $i$ in its tree. In particular, $\textsf{lca}_{0}(x,x)$ holds if and only if $x$ is a root. The predicates $\textsf{lca}_{i}$ are called lca-predicates. Note that, as opposed to the label predicates for which it may be the case that $\textsf{lab}_{\lambda}(x)$ and $\textsf{lab}_{\lambda^{\prime}}(x)$ both hold for two labels $\lambda\neq\lambda^{\prime}$, we have that, for every $x$ and $y$, $\textsf{lca}_{i}(x,y)$ holds for exactly one index $i\in\{-1,0,1,\dots,d\}$.
@@
Lemma 1.
Every formula $\varphi\in\text{FO}[\Lambda]$ to be evaluated in $\mathcal{F}_{d}[\Lambda]$ can be lca-expressed by replacing every equality predicate $x=y$ by
$\bigvee_{i\in[0,d]}\Big{(}\textsf{lca}_{i}(x,x)\land\textsf{lca}_{i}(x,y)\land\textsf{lca}_{i}(y,y)\Big{)},$
@@
$\bigvee_{i\in[0,d]}\Big{(}\textsf{lca}_{i}(x,y)\land\big{(}(\textsf{lca}_{i}(x,x)\lor\textsf{lca}_{i+1}(y,y))\lor(\textsf{lca}_{i+1}(x,x)\land\textsf{lca}_{i}(y,y))\big{)}\Big{)}.$
Given a fixed set $\Lambda$ of labels, a formula $\varphi\in\text{FO}[\Lambda]$ is called lca-reduced if (1) it is lca-expressed, and (2) it is quantifier-free. In other words, an lca-reduced formula is a Boolean combination of label- and lca predicates on free variables. Let $\overline{x}=(x_{1},\dots,x_{k})$ be a $k$-tuple of variables. For any two functions $\gamma:[k]\rightarrow 2^{\Lambda}$, and $\delta:[k]\times[k]\rightarrow[-1,d]$, let us define the formula
$\textsf{type}_{\gamma,\delta}(\overline{x})\in\text{FO}[\Lambda]$
@@
$\forall i\in\{1,\dots,k\},\;\bigwedge_{\lambda\in\gamma(i)}\textsf{lab}_{\lambda}(x_{i})\wedge\bigwedge_{\lambda\in\Lambda\smallsetminus\gamma(i)}\neg\textsf{lab}_{\lambda}(x_{i})$
@@
$\forall(i,j)\in\{1,\dots,k\}^{2},\;\textsf{lca}_{\delta(i,j)}(x_{i},x_{j}).$
We call such predicate lca-type, and we denote by $\textsf{Type}(k,d,\Lambda)$ the set of all lca-type predicates with $k$ variables. Observe that the size of $\textsf{Type}(k,d,\Lambda)$ is at most $2^{k|\Lambda|}(d+2)^{k^{2}}$, and that it can be computed from $k,d$ and $\Lambda$. Note also that, given $\gamma$ and $\delta$, it may be the case that the formula $\textsf{type}_{\gamma,\delta}(\overline{x})$ is always false. This is typically the case when the function $\delta$ yields a conjunction of incompatible lca predicates (e.g., $\delta(i,j)=5$, $\delta(i,i)=2$, and $\delta(j,j)=1$, or merely any function $\delta$ such that $\delta(i,j)\neq\delta(j,i)$). In such cases, i.e., when $\textsf{type}_{\gamma,\delta}(\overline{x})$ is always false, we say that the formula is trivial.
For any non-trivial lca-type formulas $\varphi(\overline{x})=\textsf{type}_{\gamma,\delta}(\overline{x})$, we have that $\delta$ and $\gamma$ induce a labeled forest $(F_{\varphi},L_{\varphi})\in\mathcal{F}_{d}[\Lambda]$, as follows. Each $i\in[k]$ is identified with a node $u_{i}$ in $F_{\varphi}$, and each leaf of $F_{\varphi}$ belongs to the $k$-tuple $\overline{u}_{\varphi}=(u_{1},\dots,u_{k})$. Furthermore, for every $(i,j)\in[k]\times[k]$, $\textsf{lca}_{\delta(i,j)}(u_{i},u_{j})$ holds in $F_{\varphi}$. Finally, for every $i\in[k]$, $L_{\varphi}(u_{i})=\gamma(i)$, and $L_{\varphi}(v)=\emptyset$ whenever $v\notin\overline{u}_{\varphi}$.
For any labeled forest $(F,L)$, any integer $k$, and any set of vertices $\overline{v}\in V(F)^{k}$, we denote by $F[\overline{v}]$ the subgraph of $F$ induced by the nodes in $\overline{v}$, and by all their ancestors.
Lemma 2.
Let $(F,L)\in\mathcal{F}_{d}[\Lambda]$ and let $\varphi=\textsf{type}_{\gamma,\delta}$ be a non-trivial lca-type formula with $k$ free variables. For every $\overline{v}\in V(F)^{k}$, $(F,L)\models\varphi(\overline{v})$ if and only if there exists a graph isomorphism $f$ between $F[\overline{v}]$ and $F_{\varphi}$ such that, for every $i\in[k]$, $L_{\varphi}(f(v_{i}))=L(v_{i})$.
@@
Recall that $\overline{u}_{\varphi}=(u_{1},\dots,u_{k})$ includes the leaves of $F_{\varphi}$. Let $\overline{v}=(v_{1},\dots,v_{k})$ be a set of $k$ vertices of $F$.
Suppose first that $(F,L)\models\varphi(\overline{v})$. Then, for every $(i,j)\in[k]\times[k]$, $\gamma(i)=L(v_{i})$ and $\textsf{lca}(v_{i},v_{j})=\delta(i,j)$. Let $f$ be the function that, for every $i\in[k]$, and for every $0<r<\delta(i,i)$, $f(v_{i})=u_{i}$, and $f$ maps the ancestor at distance $r$ of $v_{i}$ in $F$ to the corresponding ancestor of $u_{i}$ in $F_{\varphi}$. By construction $f$ is an isomorphism between $F[\overline{v}]$ and $F_{\varphi}$. Moreover, for every $i\in[k]$ we do have $L_{\varphi}(f(v_{i}))=L_{\varphi}(u_{i})=\gamma(i)=L(v_{i})$, as claimed.
Conversely, let us assume that there exists a graph isomorphism $f$ between $F[\overline{v}]$ and $F_{\varphi}$ such that, for every $i\in[k]$, $L_{\varphi}(f(v_{i}))=L(v_{i})$. Then we directly infer that, for every $(i,j)\in[k]\times[k]$, $\gamma(i)=L(v_{i})$ and $\textsf{lca}(v_{i},v_{j})=\delta(i,j)$. This implies that $(F,L)\models\varphi(\overline{v})$, as claimed. ∎
@@
Lemma 3.
For every lca-reduced formula $\varphi(\overline{x})\in\text{FO}[\Lambda]$, there exists a set $I\subseteq\textsf{Type}(|\overline{x}|,d,\Lambda)$, and a formula
@@
such that, for every $(F,L)\in\mathcal{F}_{d}[\Lambda]$, $\textsf{true}(\varphi,F,L)=\textsf{true}(\varphi^{\prime},F,L)$. Moreover, $\varphi^{\prime}(\overline{x})$ can be computed given only $d$, $\Lambda$, and $\varphi$. The formula $\varphi^{\prime}(\overline{x})$ is called the basic normal form of $\varphi$.
@@
Let $\varphi(\overline{x})$ be an lca-reduced formula with $k$ free variables, and let $(F,L)\in\mathcal{F}_{d}[\Lambda]$. Let $\overline{v}=(v_{1},\dots,v_{k})$ be a tuple of nodes such that $(F,L)\models\varphi(\overline{v})$. Let us define a function
$\delta_{\overline{v}}:[k]\times[k]\rightarrow[-1,d]$
such that $\delta_{\overline{v}}(i,j)=h$ if the path $P_{i}$ from $v_{i}$ to the root of the tree of $F[\overline{v}]$ containing $v_{i}$, and the path $P_{j}$ from $v_{j}$ to the root of the tree of $F[\overline{v}]$ containing $v_{j}$ share exactly $h+1$ nodes. We also define the function $\gamma_{\overline{v}}$ by
$\gamma_{\overline{v}}(i)=L(v_{i})$

### v2

@@ 2.4 Graphs of Bounded Expansion
It has been recently shown in [25] that decision, counting and optimization of FO formulas can be performed in $\mathcal{O}(1)$ rounds in CONGEST, for connected labeled graphs of bounded treedepth.
Proposition 2.7 ([25]).
@@
Let $\mathcal{G}[\Lambda]$ be a class of connected labeled graphs of bounded treedepth, and let $\varphi(x_{1},\dots,x_{k})\in FO[\Lambda]$. Then, for any and any $(G,\ell)\in\mathcal{G}[\Lambda]$, the following problems can be solved by a distributed algorithm running in $\mathcal{O}(1)$ rounds in CONGEST:
- Deciding $\varphi$:
@@
deciding whether it exists a $k$-uple of vertices $(v_{1},\dots,v_{k})$ such that$(G,\ell)\models\varphi(v_{1},\dots,v_{k})$.
- Counting the solutions of $\varphi$:
@@
computing the number of $k$-uples $(v_{1},\dots,v_{k})$ such that$(G,\ell)\models\varphi(v_{1},\dots,v_{k})$.
- Optimizing $\varphi$:
[…]
@@
Graphs of bounded expansion [54] can be defined in term of local minors, and degeneracy, as pointed out in the Introduction. It is however more convenient for us to use the following characterization. A class $\mathcal{G}$ of graphs has bounded expansion if there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and for every $k\in\mathbb{N}$, there exists a coloring of the vertices of $G$ with colors in $[f(k)]=\{1,\dots,f(k)\}$ such that, for every set $U\subseteq[f(k)]$ of cardinality at most $k$ the subgraph of $G$ induced by the nodes colored by a color in $U$ has treedepth at most $|U|$. In particular, each subset of at most $k$ colors induces a subgraph of treedepth at most $k$. For every positive integers $p$ and $k$, a $(p,k)$-treedepth coloring of a graph $G$ is a $k$-coloring of $G$ satisfying that each set $U\subseteq[k]$ of at most $p$ colors induces a graph of treedepth at most $|U|$.
Definition 2.8.
Given a function $f:\mathbb{N}\rightarrow\mathbb{N}$, a class of graphs $\mathcal{G}$ has expansion $f$ if, for every $p\in\mathbb{N}_{>0}$, every $G\in\mathcal{G}$ has a $(p,f(p))$-treedepth coloring. A class of graphs has bounded expansion if it has expansion $f$ for some function $f:\mathbb{N}\rightarrow\mathbb{N}$.
@@
A class of graphs has effective bounded expansion if it has expansion $f$ for some Turing-computable function $f$. In the following, we always assume that the considered functions $f$ are Turing-computable. Interestingly, treedepth colorings can be efficiently computed in the CONGEST model.
Proposition 2.9 ([55]).
@@
Let $\mathcal{G}$ be a class of graphs of bounded expansion $f$. There exists a function $g:\mathbb{N}\to\mathbb{N}$ such that, for any graph $G\in\mathcal{G}$, and any constant positive integer $p$, a $(p,g(p))$-treedepth coloring of $G$ can be computed in $\mathcal{O}(\log n)$ rounds in the CONGEST model.
The algorithm of [55] takes as input the function $f$, and an integer $p$, and outputs a $(p,g(p))$-treedepth coloring of the actual graph $G$ in a distributed manner. The function $g$ may however be different from the given bound $f$ on the expansion of the class $\mathcal{G}$, which may itself be different from the “best” expansion function for the class $\mathcal{G}$.
3 Distributed Model Checking of Local Formulas
We have now all the ingredients to solve the open problem in [55], as established in Theorem 1.1 restated below.
See 1.1
Our algorithm runs on labeled graphs $(G,\ell)\in\mathcal{G}[\Lambda]$. Initially, every node $v\in V$ of an $n$-node labeled graph $(G,\ell)\in\mathcal{G}[\Lambda]$ is only given its identifier $\textsf{id}(v)$ on $\mathcal{O}(\log n)$ bits, which is unique in the graph, and its labels $\ell(u)\in 2^{\Lambda}$ (in addition to a polynomial upper bound on the number $n$ of vertices, as specified in the CONGEST model). At the end of the execution of our algorithm, every node $v$ outputs $\top$ if $(G,\ell)\models\varphi(v)$, and $\bot$ otherwise.
@@ 3 Distributed Model Checking of Local Formulas
We prove Theorem 1.1 by adapting the (centralized) quantifier elimination technique of [59] (see also [18, 38]) to the distributed setting. Let $(G,\ell)$ be a labeled graph with $G\in\mathcal{G}$ and $\ell:V\to 2^{\Lambda}$, on which we aim at checking a local formula $\varphi(x)$ at every node. Quantifier elimination proceeds by induction on the depth of the formula. The formula considered at any stage of the induction may thus have several free variables $\overline{x}$ with $x_{1}=x$. We therefore describe a distributed algorithm that, in $\mathcal{O}(\log n)$ rounds, transforms both the formula $\varphi(\overline{x})$ and the node-labeling $\ell$ into a new formula $\widehat{\varphi}(\overline{x})$ and a new node labeling $\widehat{\ell}$ such that:
@@
  • $\widehat{\varphi}(\overline{x})$ is quantifier-free, but uses new unary and binary predicates,
  • $\widehat{\ell}$ uses a new set of labels $\widehat{\Lambda}\supseteq\Lambda$ for the nodes of $G$, using additional labels involved in the new unary and binary predicates, and
  • the transformation preserves locality, and, for any tuple $\overline{v}$ of $|\overline{x}|$ vertices of $G$, we have:
    $(G,\ell)\models\varphi(\overline{v})\iff(G,\widehat{\ell})\models\widehat{\varphi}(\overline{v}).$
Note that, while the transformation preserves locality, it might be the case that $\varphi$ is $r$-local for some $r$, where as $\widehat{\varphi}$ is $r^{\prime}$-local for some $r^{\prime}\geq r$. Quantifiers are eliminated one by one, and the main difficulty consists in the elimination of the existential quantifiers. As in [59], the proof works in three steps.
  • First, the transformation from $(\varphi,\ell)$ to $(\widehat{\varphi},\widehat{\ell})$ is performed on rooted forests of constant depth (cf. Subsection 3.1). This is the most technical part of the quantifier elimination process.
  • Second, the quantifier elimination is performed on graphs of bounded treedepth (cf. Subsection 3.2), by reducing the analysis of graphs of bounded treedepth to the analysis of rooted forests. This is where the set of labels is enriched.
  • Third, the analysis of graphs of bounded expansion is reduced to the analysis of graphs of bounded treedepth thanks to a $(p,f(p))$-treedepth coloring (cf. Subsection 3.3), for an appropriate $p$. This is the steps for which the preservation of the locality property needs particular care.
@@
We emphasis that, at each step of the transformation of the original formula $\varphi$, and the original labeled graph $(G,\ell)$, the computations must be efficiently performed in the CONGEST model. Typically, one needs to make sure that adding new labels to the graph can be done efficiently in CONGEST. The rooted forests involved in Subsection 3.1, as well as the graphs of bounded treedepth involved in Subsection 3.2, are subgraphs of the communication graph $G$, and thus all communications passing through the edges of the former can be implemented along the edges of the latter.
3.1 Rooted Forests of Bounded Depth
@@
In this section we adapt the quantifier elimination procedure from [59] for rooted forests of constant depth (i.e., every tree in the forest is rooted and has constant depth) to the case of local formulas, and show how to implement this procedure in CONGEST. The basic idea for eliminating the quantifiers is to encode information about the satisfiability of the formula in node-labels. This approach does increase the size of the label set. Nevertheless, by taking advantage of the simple structure of FO formulas on forests of constant depth, it is possible to bound the number of labels by a constant. This constant depends solely on the size of the formula, and on the depth of the trees.
As the locality property focuses on distances between nodes, the diameter of each tree plays an important role. Therefore, instead of considering forests with depth at most $d$ for some $d$, we consider forests in which each tree has diameter at most $d-1$ for some $d$. For every $d\in\mathbb{N}$, let us denote by $\mathcal{F}_{d}$ the set of forests in which each tree has diameter at most $d-1$. Note that, in particular, all forests of $\mathcal{F}_{d}$ have depth at most $d$. For a set of labels $\Lambda$, let us denote by $\mathcal{F}_{d}[\Lambda]$ the set of all $\Lambda$-labeled graphs $(G,\ell)$ such that $G\in\mathcal{F}_{d}$.
3.1.1 First-Order Formulas on Forests of Bounded Depth
@@
Let $\Lambda$ be a (finite) set of labels, and let $d$ be a positive integer. It is convenient to consider a specific way of encoding formulas, suitable for $\mathcal{F}_{d}[\Lambda]$. For each $i\in\{-1,0,1,\dots,d-1\}$, we define the predicate $\textsf{lca}_{i}(x,y)$ (for least common ancestor) as the predicate with free variables $x$ and $y$ that holds if the path $P_{x}$ from $x$ to the root of the tree containing $x$, and the path $P_{y}$ from $y$ to the root of the tree containing $y$ share exactly $i+1$ nodes. Moreover, we set $\textsf{lca}_{-1}(x,y)$ to be true whenever $x$ and $y$ are in different trees. Note that $\textsf{lca}_{i}(x,x)$ holds if and only if $x$ is a node at depth $i$ in its tree. In particular, $\textsf{lca}_{0}(x,x)$ holds if and only if $x$ is a root. The predicates $\textsf{lca}_{i}$ are called lca-predicates. As opposed to the label predicates, for which it may be the case that $\textsf{lab}_{\lambda}(x)$ and $\textsf{lab}_{\lambda^{\prime}}(x)$ both hold for two different labels $\lambda\neq\lambda^{\prime}$, we have that, for every $x$ and $y$, $\textsf{lca}_{i}(x,y)$ holds for exactly one index $i\in\{-1,0,1,\dots,d-1\}$.
@@
Lemma 3.1.
Every formula $\varphi\in\mathrm{FO}[\Lambda]$ to be evaluated in $\mathcal{F}_{d}[\Lambda]$ can be lca-expressed by replacing every equality predicate $x=y$ by
$\bigvee_{i\in[0,d-1]}\Big{(}\textsf{lca}_{i}(x,x)\land\textsf{lca}_{i}(x,y)\land\textsf{lca}_{i}(y,y)\Big{)},$
@@
$\bigvee_{i\in[0,d-2]}\Big{(}\textsf{lca}_{i}(x,y)\land\big{(}(\textsf{lca}_{i}(x,x)\land\textsf{lca}_{i+1}(y,y))\lor(\textsf{lca}_{i+1}(x,x)\land\textsf{lca}_{i}(y,y))\big{)}\Big{)}.$
Given a fixed set $\Lambda$ of labels, a formula $\varphi\in\mathrm{FO}[\Lambda]$ is called lca-reduced if (1) it is lca-expressed, and (2) it is quantifier-free. In other words, an lca-reduced formula is a Boolean combination of label and lca predicates on free variables.
Definition 3.2.
Let $\overline{x}=(x_{1},\dots,x_{k})$ be a $k$-tuple of variables. For any two functions $\gamma:[k]\rightarrow 2^{\Lambda}$, and $\delta:[k]\times[k]\rightarrow[-1,d-1]$, let us define the formula
$\textsf{type}_{\gamma,\delta}(\overline{x})\in\mathrm{FO}[\Lambda]$
@@
[…]
@@
[…]
We call such predicate lca-type, and we denote by $\textsf{Type}(k,d,\Lambda)$ the set of all lca-type predicates with $k$ variables.
Observe that the size of $\textsf{Type}(k,d,\Lambda)$ is at most $2^{k|\Lambda|}(d+1)^{k^{2}}$, and that it can be computed from $k,d$ and $\Lambda$. Note also that, given $\gamma$ and $\delta$, it may be the case that the formula $\textsf{type}_{\gamma,\delta}(\overline{x})$ is always false. This is typically the case when the function $\delta$ yields a conjunction of incompatible lca predicates (e.g., $\delta(i,j)=5$, $\delta(i,i)=2$, and $\delta(j,j)=1$, or merely any function $\delta$ such that $\delta(i,j)\neq\delta(j,i)$). In such cases, i.e., when $\textsf{type}_{\gamma,\delta}(\overline{x})$ is always false, we say that the formula is trivial.
For any non-trivial lca-type formulas $\varphi(\overline{x})=\textsf{type}_{\gamma,\delta}(\overline{x})$, we have that $\delta$ and $\gamma$ induce a labeled forest $(F_{\varphi},\ell_{\varphi})\in\mathcal{F}_{d}[\Lambda]$, as follows. Each $i\in[k]$ is identified with a node $u_{i}$ in $F_{\varphi}$, and each leaf of $F_{\varphi}$ belongs to the $k$-tuple $\overline{u}_{\varphi}=(u_{1},\dots,u_{k})$. Furthermore, for every $(i,j)\in[k]\times[k]$, $\textsf{lca}_{\delta(i,j)}(u_{i},u_{j})$ holds in $F_{\varphi}$. Finally, for every $i\in[k]$, $\ell_{\varphi}(u_{i})=\gamma(i)$, and $\ell_{\varphi}(v)=\emptyset$ whenever $v\notin\overline{u}_{\varphi}$.
For any labeled forest $(F,\ell)$ with $F=(V,E)$, any integer $k$, and any set of vertices $\overline{v}\in V^{k}$, let us denote by $F[\overline{v}]$ the subgraph of $F$ induced by all the nodes in $\overline{v}$, and their ancestors.
Lemma 3.3.
Let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$ with $F=(V,E)$, and let $\varphi(\overline{x})=\textsf{type}_{\gamma,\delta}(\overline{x})$ be a non-trivial lca-type formula with $k$ free variables. For every $\overline{v}=(v_{1},\dots,v_{k})\in V^{k}$, $(F,\ell)\models\varphi(\overline{v})$ if and only if there exists a graph isomorphism $f$ between $F[\overline{v}]$ and $F_{\varphi}$ such that, for every $i\in[k]$, $\ell_{\varphi}(f(v_{i}))=\ell(v_{i})$.
@@
Recall that $\overline{u}_{\varphi}=(u_{1},\dots,u_{k})$ includes the leaves of $F_{\varphi}$. Let $\overline{v}=(v_{1},\dots,v_{k})$ be a set of $k$ vertices of $F$. Suppose first that $(F,\ell)\models\varphi(\overline{v})$. Then, for every $(i,j)\in[k]\times[k]$,
$\gamma(i)=\ell(v_{i}),\text{ and }\textsf{lca}(v_{i},v_{j})=\delta(i,j).$
Let $f$ be the function that, for every $i\in[k]$, $f(v_{i})=u_{i}$, and, for every $r\in\{1,\dots,\delta(i,i)-1\}$, $f$ maps the ancestor of $v_{i}$ in $F$ at distance $r$ from $v_{i}$ to the corresponding ancestor of $u_{i}$ in $F_{\varphi}$. By construction $f$ is an isomorphism between $F[\overline{v}]$ and $F_{\varphi}$. Moreover, for every $i\in[k]$, we do have $\ell_{\varphi}(f(v_{i}))=\ell_{\varphi}(u_{i})=\gamma(i)=\ell(v_{i})$, as claimed.
Conversely, let us assume that there exists a graph isomorphism $f$ between $F[\overline{v}]$ and $F_{\varphi}$ such that, for every $i\in[k]$, $\ell_{\varphi}(f(v_{i}))=\ell(v_{i})$. Then we directly infer that, for every $(i,j)\in[k]\times[k]$, $\gamma(i)=\ell(v_{i})$ and $\textsf{lca}(v_{i},v_{j})=\delta(i,j)$. This implies that $(F,\ell)\models\varphi(\overline{v})$, as claimed. ∎
@@
Lemma 3.4.
For every lca-reduced formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$, there exists a set $I\subseteq\textsf{Type}(|\overline{x}|,d,\Lambda)$, and a formula
@@
such that, for every $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$, $\textsf{true}(\varphi,F,\ell)=\textsf{true}(\varphi^{\prime},F,\ell)$. Moreover, $\varphi^{\prime}(\overline{x})$ can be computed given only $d$, $\Lambda$, and $\varphi$. The formula $\varphi^{\prime}(\overline{x})$ is called the basic normal form of $\varphi$.
@@
Let $\varphi(\overline{x})$ be an lca-reduced formula with $k$ free variables, and let $(F,\ell)\in\mathcal{F}_{d}[\Lambda]$. Let $\overline{v}=(v_{1},\dots,v_{k})$ be a tuple of nodes such that $(F,\ell)\models\varphi(\overline{v})$. Let us define the function
$\delta_{\overline{v}}:[k]\times[k]\rightarrow[-1,d-1]$
by $\delta_{\overline{v}}(i,j)=h$ if the path from $v_{i}$ to the root of the tree of $F[\overline{v}]$ containing $v_{i}$, and the path from $v_{j}$ to the root of the tree of $F[\overline{v}]$ containing $v_{j}$ share exactly $h+1$ nodes. We also define the function $\gamma_{\overline{v}}$ by
$\gamma_{\overline{v}}(i)=\ell(v_{i})$